OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 A. 1 B. 4 C. 5 D. 25

OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 2 points) Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution Question 4 (Multiple Choice Worth 1 points) Solve for x: 3x − 5 = 2x + 6 1 −1 11 −11 Question 5 (Essay Worth 4 points) Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

@zordoloom PLEASE assist me with these

OpenStudy (anonymous):

Lol, May I ask what class this is for? BTW the answer for the fist one is x=1. OR A

OpenStudy (anonymous):

@zordoloom thank you for all your help you have been giving me, i mean it. and Algerbra 1

OpenStudy (anonymous):

oh, How I miss algebra 1. Question 3, No solution.

OpenStudy (anonymous):

haha and thanks

OpenStudy (anonymous):

Question 4, x=11, Or the third one.

OpenStudy (anonymous):

For the fifth one, Just make up any equation, and set it equal to itself.

OpenStudy (anonymous):

so what does equal to itself mean??

OpenStudy (anonymous):

yo you are so fing smart bro, like whats your highest math you ever took??

OpenStudy (anonymous):

So for example 5x+7=5x+7

OpenStudy (anonymous):

I'm taking differential equation, which is past Calculus.

OpenStudy (anonymous):

5x=5x.

OpenStudy (anonymous):

This would give you an all real solutions.

OpenStudy (anonymous):

dag dude your so smart haha, can you start helping me everyday or when ur not busy?? and when im doing the two-step equation with All Real Numbers as the solution. how would i Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

So, 5x+7=5x+7 Subtract the 7's They pretty much cancel out 5x=5x Now divide by 5 x=x All real solutions.

OpenStudy (anonymous):

Well, whenever you see me logged on, just message me if you need help.

OpenStudy (anonymous):

Okay and thank you so much dude, can u help me some more??

OpenStudy (anonymous):

Just make a new post.

OpenStudy (anonymous):

I have like 14 more questions to do, could u help me finish those and that will be it for tonight??

OpenStudy (anonymous):

@zordoloom ^^^

OpenStudy (anonymous):

I can help you, but what good does it do for you? Its up to you.

OpenStudy (anonymous):

Okay ill make a new post